# Diffusion Model Patching via Mixture-of-Prompts

### v1

@@ 4.2 Design Choices
Prompt balancing loss acts as a soft constraint for the gating function, helping to mitigate biased selections of prompts when producing a mixture-of-prompts. We study the effects of two types of balancing losses: importance balancing loss and load balancing loss. As shown in Table 4(e), using both types of losses enables the diffusion model to reach its peak performance. When used individually, each type of oss contributes slightly to performance improvement, but when used together, they exhibit complementary effects, boosting overall performance. This indicates that balancing both the weight of the activated prompts and the number of activated prompts across different timesteps is crucial for creating an effective mixture-of-prompts. Consequently, we use both importance balancing loss and load balancing loss for prompt balancing.
Figure 5: Prompt activation. Brighter means stronger activation.
@@
Table 5: Gating condition.
4.3 Analysis
(a) Unconditional Image Generation on FFHQ [25]
Baseline
+ PT
+ DMP
(b) Text-to-Image Generation on MS-COCO [33].
Baseline
+ PT
+ DMP
Figure 6: Qualitative comparison among the baseline (DiT-B/2 [43]), naive prompt tuning (PT) applied to the baseline [29], and DMP applied to the baseline on (a) FFHQ [25] and (b) MS-COCO [33] datasets.
Prompt activation.
The gating function plays a pivotal role in dynamically crafting mixtures-of-prompts from a set of learnable prompts, based on the noise level present in the input. This is depicted in Fig. 5, where the activation is visually highlighted using colors. As the denoising process progresses, the selection of prompts exhibits significant variation across different timesteps. At higher timesteps, where noise levels are greater, the gating function tends to utilize a broader array of prompts. Conversely, at lower timesteps, as the noise diminishes, the prompts become more specialized, focusing narrowly on specific features of the input that demand closer attention. This strategic deployment of prompts allows the model to form specialized "experts" at each denoising step. Such adaptability enables the model to adjust its approach based on the current noise level, effectively leveraging different prompts to enhance its overall performance. Using mixture-of-prompts tailored to different noise levels ensures that the image generation process is both flexible and precise, catering to the specific needs dictated by the input’s noise characteristics at each step.
Gating condition.
In class-conditional image generation and text-to-image generation tasks, conditional guidance plays a crucial role in determining the outcome of generated images [8, 18]. To study the impact of conditional guidance on selecting mixtures-of-prompts, we evaluate the performances of two cases: one where the gating function $\mathcal{G}$ receives only the timestep embedding ${\bm{t}}$, and the other where it receives both ${\bm{t}}$ and the class or text condition embedding ${\bm{c}}$. In the latter case, we modify the input of the gating function in Eq. 5, from ${\bm{t}}$ to ${\bm{t}}+{\bm{c}}$. Our analysis, presented in Section 4.2, indicates that on the ImageNet dataset, both methods perform equivalently, whereas on the COCO dataset, using ${\bm{t}}+{\bm{c}}$ yields superior performance compared to using ${\bm{t}}$ alone. This suggests that incorporating conditional guidance can help in determining how to combine prompts at each denoising step to generate an image that aligns well with the text condition. Consequently, we use ${\bm{t}}+{\bm{c}}$ as the default input for the gating function in the text-to-image task.
Qualitative analysis.
Fig. 6 presents a visual comparison between three methods: the baseline DiT model, prompt tuning [29], and our DMP. These methods are evaluated on unconditional, text-to-image generation tasks using FFHQ [25] and COCO [33], respectively. Overall, DMP demonstrates realistic and natural images with fewer artifacts. We provide additional qualitative results in Appendix E.
5 Conclusion
In this study, we introduced Diffusion Model Patching (DMP), a lightweight and computationally efficient method for enhancing pre-trained diffusion models that have already converged. Most pre-trained diffusion models do not explicitly integrate denoising-stage-specificity into their model design. Given that the characteristics of denoising stages differ slightly, we aim to modify the model to be aware of such differences. The key idea is to integrate timestep-specific learnable prompts to patch the model’s behavior at every timestep. To achieve this, DMP leverages dynamic gating, which adaptively combines prompts based on the current timestep (or noise level) during the denoising process. This allows our method to scale effectively to thousands of denoising steps. Our results demonstrate that while fine-tuning the model on the original dataset does not lead to further improvement, DMP uniquely enhances performance. When applied to the DiT-XL backbone, DMP yields substantial FID gain of 10.38% in 50K iterations, with just 1.43% extra parameters on FFHQ 256$\times$256. Notably, DMP’s effectiveness extends across different model sizes and various image generation tasks.
Limitations. Our DMP method adopts a prompt-adding strategy to ensure stable training and maintain sampling speed. However, since the number of input patches is fixed, the flexibility in the number of prompts is limited. Extending our DMP with a prepend approach while maintaining stable training is an interesting future direction.
Appendix
Appendix A Additional Related Work
Prompt-based learning.
Recent progress in NLP has shifted towards leveraging pre-trained language models (LMs) using textual prompts [44, 46, 52, 34] to guide models to perform target tasks or produce desired outputs without additional task-specific training. With strategically designed prompts, models like GPT-3 [3] have shown impressive generalization across various downstream tasks, even under few-shot or zero-shot conditions. Prompt tuning [31, 29, 36, 35] treats prompts as learnable parameters optimized with supervision signals from downstream training samples while keeping the LM’s parameters fixed. Similar principles have also been explored in visual [22, 1] and vision-and-language [71, 70] domains. To the best of our knowledge, there is currently no direct extension of prompt tuning in enhancing the in-domain performance of diffusion models. While Prompt2prompt [15] and Textual Inversion [11] share similar properties with prompt tuning, their focus is on customized editing or personalized content generation. In this work, we propose to leverage prompt tuning to enhance the stage-specific capabilities of diffusion models. With only a small number of prompts, we can effectively scale to thousands of denoising steps via a mixture-of-prompts strategy.
Appendix B Preliminary
Diffusion models.
Diffusion models [8, 55] generate data by reversing a pre-defined diffusion process (or forward process), which sequentially corrupts the original data ${\bm{x}}_{0}$ into noise over a series of steps $t\in\{1,\dots,T\}$.
where $10<\beta_{t}<$ 1 is a variance schedule controlling the amount of noise added at each step. This process results in data that resembles pure noise $\mathcal{N}(0,\mathbf{I})$ at step $T$ (often $T=1000$). The reverse process aims to reconstruct the original data by denoising, starting from noise and moving backward to the initial state ${\bm{x}}_{0}$. This is modeled by a neural network parameterized by $\bm{\theta}$ that learns the conditional distribution $p_{\bm{\theta}}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. The network is trained by optimizing a weighted sum [17] of denoising score matching losses [60] over multiple noise scales [56]. In practice, the network predicts the noise $\bm{\epsilon}$ added at each forward step, rather than directly predicting ${\bm{x}}_{t-1}$ from ${\bm{x}}_{t}$, using the objective function:
Appendix C Prompt Balancing Loss
During the training stage of our DMP method, we adopt a prompt balancing loss to prevent the gate from selecting only a few specific prompts, a problem known as mode collapse. The prompt balancing loss is inspired by the balancing loss used in mixture-of-experts. Our prompt balancing loss includes the importance loss, which ensures prompts are selected uniformly within a batch, and the load balancing loss, which prevents the existence of unselected prompts in a batch. When defining the $n$-th prompt gating in the $i$-th DiT-block layer as $g^{i}_{n}$, the formulations of the importance loss and load balancing loss are as follows:
$\begin{split}\mu_{i}=\frac{1}{N}\sum^{N-1}_{n=0}{g^{i}_{n}},&\;\;\;\;\;\sigma_{i}=\frac{1}{N}\sum^{N-1}_{n=0}{(g^{i}_{n}-\mu_{i})^{2}}\end{split}$ (9)
$\begin{split}L_{importance}=\sum^{L-1}_{i=0}{\frac{\sigma_{i}}{\mu_{i}^{2}+\epsilon}},&\;\;\;\;\;L_{Load}=\frac{1}{L}\sum^{L-1}_{i=0}{I(g^{i}_{n}<0)},\end{split}$ (10)
where $L$ is the total number of DiT block layers, $N$ is the total number of prompts, $\epsilon$ is $51e^{-}$ 5 to prevent division by zero, and $I$ is the indicator function that counts the number of elements satisfying the condition $(g^{i}_{n}>0)$. The importance loss uses the squared coefficient of variation, which makes the variance value robust against the mean value. With these prompt balancing losses, we regularize the prompt selection in the gate, ensuring that there are few unselected prompts, as shown in Fig. 5 of main manuscript.
Appendix D Implementation Details
Our experiments followed this setup for further training of pretrained DiT-B/L/XL models [43]. Firstly, we used a diffusion timestep T of 1000 for training and DDPM with 250 steps [17] for sampling. For beta scheduling, cosine scheduling [40] was used for the FFHQ [25] and MS-COCO datasets [33], while linear scheduling was used for the ImageNet dataset [7]. For text-to-image generation (MS-COCO) and class-conditional image generation (ImageNet) tasks, we adopted classifier-free guidance [18] with a guidance scale of 1.5. The batch size was set to 128, and random horizontal flipping was applied to the training data. We used the AdamW optimizer [38] with a fixed learning rate of 1e-4 and no weight decay. Originally, the exponential moving average (EMA) technique is utilized for stable training of DiT models. However, since our method involves further training on pretrained models with only a few training steps, we did not adopt the EMA strategy. All experiments were conducted using a single NVIDIA A100 GPU.
Appendix E Additional Qualitative Results on ImageNet
@@
Figures 7, 8, 9 and 10 illustrates the generated images by pre-trained DiT-XL with DMP trained on 50K iterations. The results demonstrate that highly realistic images can be generated, even with relatively limited training.
Figure 7: Uncurated 256×256 DiT-XL/2+DTR samples.
@@
Figure 8: Uncurated 256×256 DiT-XL/2+DTR samples.
@@
Figure 9: Uncurated 256×256 DiT-XL/2+DTR samples.
@@
Figure 10: Uncurated 256×256 DiT-XL/2+DTR samples.

### v3

@@ Introduction
Prompt balancing. Prompt balancing loss acts as a soft constraint for the gating function, mitigating biased selections of prompts when producing a mixture-of-prompts. We study the impact of two types of balancing losses by altering the coefficient values for load balancing loss and importance balancing loss. As shown in Table 4(e), using both types of losses equally enables the diffusion model to reach its peak performance. This indicates that balancing both the number and weight of the activated prompts across different timesteps is crucial for creating an effective mixture-of-prompts. Consequently, we employ equal proportions of importance balancing and load balancing losses for prompt balancing.
Analysis
Prompt activation. The gating function plays a pivotal role in dynamically crafting mixtures-of-prompts from a set of learnable prompts, based on the noise level present in the input. This is depicted in Fig. 5, where the activation is visually highlighted using colors. As the denoising process progresses, the selection of prompts exhibits significant variation across different timesteps. At higher timesteps with high noise levels, the gating function tends to utilize a broader array of prompts. Conversely, at lower timesteps, as the noise diminishes, the prompts become more specialized, focusing narrowly on specific features of the input that demand closer attention. This strategic deployment of prompts allows the model to form specialized ”experts” at each denoising step, catering to the specific needs dictated by the input’s noise characteristics and enhancing the model’s performance.
Qualitative analysis. Fig. 6 presents a visual comparison between three methods: the baseline DiT model, prompt tuning (Lester, Al-Rfou, and Constant 2021), and our DMP. These methods are evaluated on unconditional, text-to-image generation tasks using FFHQ (Karras, Laine, and Aila 2019) and COCO (Lin et al. 2014), respectively. DMP generates realistic and natural images with fewer artifacts.
Additional results and analysis. Appendix provides a theoretical grounding of DMP, experiments on training diffusion models from scratch with DMP, applying DMP on DiT-XL/2, comparisons with LoRA (Hu et al. 2021), and ablations of gating conditions. It also examines the structural bias of DMP and provides additional qualitative results.
Conclusion
We introduced Diffusion Model Patching (DMP), a simple method for further enhancing pre-trained diffusion models that have already converged. By incorporating timestep-specific learnable prompts and leveraging dynamic gating, DMP adapts the model’s behavior dynamically across thousands of denoising steps. This design enables DMP to effectively address the variations inherent in denoising stages, which are often overlooked in existing diffusion model architectures. Our results demonstrate that DMP achieves significant performance gains without the need for extensive retraining. Applied to the DiT-L/2 backbone, DMP delivered a 10.38% improvement in FID after just 50K iterations, with a minimal parameter increase of 1.43% on the FFHQ 256×256 dataset. Additionally, its adaptability across different models and image generation tasks underscores its potential as a versatile enhancement method for diffusion models.
Appendix
Appendix A Theoretical Grounding of DMP
Our work is grounded in the framework of Multi-task Learning (Caruana 1997), specifically leveraging parameter sharing and parameter separation notions. We detail the theoretical framework for our approach:
1) Parameter Sharing: We build upon an already converged diffusion model, parameterized by $\theta$, which serves as the shared trunk. This shared trunk has been trained across all timesteps $t\in T$, effectively functioning as a multi-task model by learning a general representation applicable to all tasks (timesteps).
2) Parameter Separation: To enhance performance, we introduce new prompts parameterized by $\phi$. These prompts are designed to learn specialized parameters for each timestep, thereby explicitly addressing the unique aspects of each task within the multi-task learning framework.
Mathematically, let $x_{t}$ denote the input at timestep $t$, and let $f(x_{t};\theta)$ represent the output of the shared trunk. The introduction of prompts modifies this output, which can be expressed as:
$y_{t}=f(x_{t};\theta)+g(x_{t},t;\phi)$ (4)
where $g(x_{t},t;\phi)$ represents the contribution of the prompt-specific parameters. During training, the shared trunk parameters $\theta$ remain fixed, while the prompt parameters $\phi$ are optimized. This ensures that the learning process focuses on fine-tuning the model for each specific timestep without altering the foundational multi-task model. The gate mechanism dynamically adjusts the weight of each prompt’s contribution based on the timestep, effectively performing parameter separation. This can be represented as:
$\alpha_{t}=\text{Gate}(t)$ (5)
$y_{t}=f(x_{t};\theta)+\alpha_{t}\cdot g(x_{t},t;\phi)$ (6)
where $\alpha_{t}$ is the weight determined by the gate for timestep $t$.
By reusing the original data used for pre-training (in-domain) to learn the prompts, we ensure that the performance improvements are due to the Multi-task learning effect rather than transfer learning.
Appendix B Extended Related Work
Prompt-based learning.
Recent progress in NLP has shifted towards leveraging pre-trained language models (LMs) using textual prompts (Petroni et al. 2019; Radford et al. 2019; Schick and Schütze 2020; Liu et al. 2023a) to guide models to perform target tasks or produce desired outputs without additional task-specific training. With strategically designed prompts, models like GPT-3 (Brown et al. 2020) have shown impressive generalization across various downstream tasks, even under few-shot or zero-shot conditions. Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Liu et al. 2023b, 2021) treats prompts as learnable parameters optimized with supervision signals from downstream training samples while keeping the LM’s parameters fixed. Similar principles have also been explored in visual (Jia et al. 2022; Bahng et al. 2022) and vision-and-language (Zhou et al. 2022b, a) domains. To the best of our knowledge, there is currently no direct extension of prompt tuning in enhancing the in-domain performance of diffusion models. While Prompt2prompt (Hertz et al. 2022) and Textual Inversion (Gal et al. 2022) share similar properties with prompt tuning, their focus is on customized editing or personalized content generation. In this work, we propose to leverage prompt tuning to enhance the stage-specific capabilities of diffusion models. With only a small number of prompts, we can effectively scale to thousands of denoising steps via a mixture-of-prompts strategy.
Appendix C Preliminaries
Diffusion models.
Diffusion models (Dhariwal and Nichol 2021; Song, Meng, and Ermon 2020) generate data by reversing a pre-defined diffusion process (or forward process), which sequentially corrupts the original data ${\bm{x}}_{0}$ into noise over a series of steps $t\in\{1,\dots,T\}$.
where $10<\beta_{t}<$ 1 is a variance schedule controlling the amount of noise added at each step. This process results in data that resembles pure noise $\mathcal{N}(0,\mathbf{I})$ at step $T$ (often $T=1000$). The reverse process aims to reconstruct the original data by denoising, starting from noise and moving backward to the initial state ${\bm{x}}_{0}$. This is modeled by a neural network parameterized by $\bm{\theta}$ that learns the conditional distribution $p_{\bm{\theta}}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. The network is trained by optimizing a weighted sum (Ho, Jain, and Abbeel 2020) of denoising score matching losses (Vincent 2011) over multiple noise scales (Song and Ermon 2019). In practice, the network predicts the noise $\bm{\epsilon}$ added at each forward step, rather than directly predicting ${\bm{x}}_{t-1}$ from ${\bm{x}}_{t}$, using the objective function:
Prompt tuning.
$\hat{{\bm{y}}}^{\prime}={{{\bm{f}}}^{\textit{learn}}_{\bm{\theta}^{\prime}}}({{\bm{x}}}),$ (9)
where $\hat{{\bm{y}}}^{\prime}$ is the refined prediction, ${{\bm{f}}}_{\bm{\theta}^{\prime}}$ is the modified model, and The superscript $learn$ indicates learnable parameters, while $frozen$ indicates frozen parameters. This process is often computationally expensive and resource-intensive, as it requires storing and updating the full model parameters. In contrast, prompt tuning aims to enhance the output $\hat{{\bm{y}}}$ by directly modifying the input ${\bm{x}}$:
$\hat{{\bm{y}}}^{\prime}={{{\bm{f}}}^{\textit{frozen}}_{\bm{\theta}}}({{\bm{x}}}^{learn}_{p}).$ (10)
Previous works (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Jia et al. 2022; Zhou et al. 2022b; Wang et al. 2022) commonly define ${{{\bm{x}}}^{learn}_{p}}=[{{{\bm{p}}}^{learn}};{{{\bm{x}}}}]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation. However, we take a different approach by directly adding prompts to the input, aiming to more explicitly influence the input itself, thus ${{{\bm{x}}}^{learn}_{p}}={{{\bm{p}}}^{learn}}+{{{\bm{x}}}}$. Prompts are optimized via gradient descent, similar to conventional fine-tuning, but without changing the model’s parameters.
Appendix D Architecture of Diffusion Models
Architecture of DiT.
The DiT model (Peebles and Xie 2022) is a diffusion model that utilizes a transformer-based (Vaswani et al. 2017) DDPM (Ho, Jain, and Abbeel 2020), operating within the latent space for image generation tasks. The architecture begins by employing a pre-trained Variational Autoencoder (VAE) (Kingma and Welling 2013) from Stable Diffusion (Rombach et al. 2022) to encode input images into latent codes of shape $H\times W\times D$. For example, an image with dimensions $256\times 256\times 3$ is encoded into a latent code of size $32\times 32\times 4$.
Noise is then added to the latent code, which is a standard training procedure in diffusion models, allowing the model to learn the process of denoising. The noisy latent code is divided into $N$ fixed-size patches, each of which is linearly embedded with shape $K\times K\times D$. Positional encodings (Vaswani et al. 2017) are added to these embedded patches, transforming them into a sequence of vectors that serve as the input tokens for the transformer model.
The core of the DiT architecture consists of a series of $L$ DiT blocks. Each block includes Multi-Head Self-Attention (Vaswani et al. 2017), Feed-Forward Networks, Layer Normalization (Ba, Kiros, and Hinton 2016), and residual connections (He et al. 2016). These components work together to allow the model to focus on different parts of the input sequence, capture complex dependencies, and stabilize the training process. The blocks are conditioned with timestep embeddings ${\bm{t}}$ , which provide the model with information about the stage of the denoising process, and can also be optionally conditioned with class or text embeddings. After the last DiT block, the noisy latent patches undergo the final Layer Normalization and are linearly decoded into a $K\times K\times 2D$ tensor ($D$ for noise prediction and another $D$ for diagonal covariance prediction). Finally, the decoded tokens are rearranged to match the original shape $H\times W\times D$.
Architecture of Stable Diffusion.
Stable Diffusion (Rombach et al. 2022) employs a UNet-based architecture (Ho, Jain, and Abbeel 2020) for its diffusion model, also operating in the latent space to achieve efficient and high-quality image generation. Similar to the DiT model, Stable Diffusion starts by using a pre-trained VAE to encode input images into latent codes. For instance, an image of size $256\times 256\times 3$ is transformed into a latent code of size $32\times 32\times 4$.
Noise is added to the latent codes according to a specific timestep sampled from the PNDM timestep scheduler (Liu et al. 2022), which is crucial for the model to learn the denoising process during training. The noisy latent codes are then processed through a series of attention blocks (Vaswani et al. 2017) within the UNet architecture. These blocks include Multi-Head Self-Attention, which allows the model to focus on important features within the latent representation, and Cross-Attention, which is particularly important in Stable Diffusion for integrating conditioning information like text or class embeddings into the image generation process. Feed-Forward Networks, Layer Normalization, and residual connections are also used within these blocks to enhance learning capacity, stabilize the training process, and maintain gradient flow.
After processing through the attention blocks, the latent representation undergoes a final Layer Normalization and is then decoded back to its original shape through a convolutional layer, effectively reconstructing the latent code with the original shape $H\times W\times D$.
Figure 7: Structural bias of prompts. Temporal bias is calculated by a patch-wise average of prompt activations over 1000 timesteps. Spatial bias is calculated by a timestep-wise average of prompt activations over 256 patches.
Appendix E Prompt Balancing Loss
During the training stage of our DMP method, we adopt a prompt balancing loss to prevent the gate from selecting only a few specific prompts, a problem known as mode collapse. The prompt balancing loss is inspired by the balancing loss used in mixture-of-experts (Shazeer et al. 2017). Our prompt balancing loss includes the load loss, which prevents the existence of unselected prompts, and the importance loss, which ensures the selected prompts have uniform weights. When defining the $n$-th prompt gating weights in the $i$-th DiT-block layer as $g^{i}_{n}$, the formulations of the load balancing loss and importance loss are as follows:
$L_{Load}=\frac{1}{L}\sum^{L-1}_{i=0}\sum^{N-1}_{n=0}{\mathbb{I}(g^{i}_{n}<0)},$ (11)
$\begin{split}&\tilde{g}_{n}=\sum^{L-1}_{i=0}{g^{i}_{n}},\\ \mu=\frac{1}{N}\sum^{N-1}_{n=0}{\tilde{g}_{n}},&\quad\sigma=\frac{1}{N}\sum^{N-1}_{n=0}{(\tilde{g}_{n}-\mu)^{2}},\end{split}$ (12)
$L_{importance}={\frac{\sigma}{\mu^{2}+\epsilon}},$ (13)
where $L$ is the total number of DiT block layers, $N$ is the total number of prompts, $\epsilon$ is $1e^{-5}$ to prevent division by zero, and $\mathbb{I}$ is the indicator function that counts the number of elements satisfying the condition $(g^{i}_{n}<0)$. The importance loss uses the squared coefficient of variation, which makes the variance value robust against the mean value. With these prompt balancing losses, we regularize the prompt selection in the gate, ensuring that there are few unselected prompts, as shown in Fig. 5 of main manuscript.
Appendix F Implementation Details
Our experiments followed this setup for further training of pretrained DiT-B/L models (Peebles and Xie 2022) and Stable Diffusion v1.5 model (Rombach et al. 2022). Firstly, we used a diffusion timestep T of 1000 for training and DDPM with 250 steps (Ho, Jain, and Abbeel 2020) for DiT models and PNDM (Liu et al. 2022) for Stable Diffusion v1.5 model during sampling. For beta scheduling, cosine scheduling (Nichol and Dhariwal 2021) was used for the FFHQ (Karras, Laine, and Aila 2019) and MS-COCO datasets (Lin et al. 2014), while linear scheduling was used for the ImageNet dataset (Deng et al. 2009) and Laion5B (Schuhmann et al. 2022). For text-to-image generation (MS-COCO) and class-conditional image generation (ImageNet) tasks, we adopted classifier-free guidance (Ho and Salimans 2022) with a guidance scale of 1.5, while for Stable Diffusion v1.5 (Laion5B), we used a guidance scale of 7.5. The batch size was set to 128, and random horizontal flipping was applied to the training data. We used the AdamW optimizer (Loshchilov and Hutter 2017) with a fixed learning rate of 1e-4 and no weight decay. Originally, the exponential moving average (EMA) technique is utilized for stable training of DiT models. However, since our method involves further training on pretrained models with only a few training steps, we did not adopt the EMA strategy. All experiments were conducted using a single NVIDIA A100 GPU.
Appendix G Structural Bias of DMP
Our DMP approach, which adds identical prompts to image tokens at corresponding positions, has the potential to introduce a structural bias. To evaluate the extent and impact of this bias, we conducted the following analyses on DiT-B/2 model (Peebles and Xie 2022) with FFHQ dataset (Karras, Laine, and Aila 2019): We averaged the attention weights across the same patch positions over different images throughout the entire temporal axis (i.e., across different timesteps). This helps us understand if the same patches in different images receive consistent attention, indicating a temporal bias. Similarly, averaging across the same timestep positions over different images allows us to evaluate if different patches within the same timestep receive consistent attention, revealing a spatial bias.
1) Temporal bias: Our analysis shows that the same patches across different images at various timesteps do exhibit some level of consistent attention. This indicates that there is a temporal structural bias introduced by the prompts.
2) Spatial bias: Similarly, within the same timestep, different patches across various images also exhibit consistent attention patterns, indicating a spatial structural bias.
Note that the strength of the structural bias varies depending on the specific configurations and the dataset used.
Visualization.
In Fig. 7, we show the prompt activation (activated by gating mechanism conditioned on the timestep) averaged across multiple samples, which might indicate a type of bias in the “input”. In our setup, all parameters are fixed except for the prompt, which is added directly to the image patch token. However, this might not fully reflect the impact on the “final image”. Fundamentally, the role of the prompts in DMP is not to learn completely new information. During fine-tuning, we train on the same data distribution as in pre-training. Therefore, the prompts activated at each timestep are meant to learn the fine-grained, timestep-specific features that the pre-trained diffusion model might have missed. Hence, it does not introduce entirely new structural patterns unseen in a well-converged diffusion model. Empirically, in our qualitative observations, we did not find any noticeable correlations or significant structural biases in the image introduced by specific prompts compared to the larger structural bias inherent in the multi-task diffusion model’s shared trunk. No regular patterns or meaningful structural biases were observed in the final images.
Statistics.
To numerically study the impacts of prompts on image-level, we conducted an extended analysis focusing on the spatial dimensions of the latent representations. The statistics are shown in Table 5. We generated 1,000 latent representations for both the baseline DiT-B/2 model and the DiT-B/2 + DMP. Each latent representation had a shape of (4, 32, 32), where 4 represents the number of channels, and 32x32 corresponds to the spatial dimensions. We computed the Pearson correlation coefficient matrix for each set of 1,000 latent representations, resulting in a (1000, 1000) correlation matrix. This matrix captures the pairwise correlation between all latent representations within the set. To focus on unique pairs, we excluded the diagonal elements (which represent self-correlation) and computed the max, min, mean, and standard deviation of the off-diagonal elements. Similarly, we normalized each latent representation vector and computed the cosine similarity matrix, also of shape (1000, 1000). Excluding the diagonal, we calculated the max, min, mean, and standard deviation for the off-diagonal elements. Notably, the mean correlation and cosine similarity values are lower in the DiT-B model with DMP compared to the baseline. The reduction in mean correlation and cosine similarity implies that the latent representations generated by the DiT-B + DMP are less correlated and less similar to each other compared to those from the baseline. This suggests that DMP may contribute to enhanced diversity in the generated samples.
Model Metric Max Min Mean Std
DiT-B/2 Correlation 0.7064 -0.4424 0.1715 0.1172
Similarity 0.7443 -0.4676 0.1764 0.1234
+DMP Correlation 0.7285 -0.4770 0.1410 0.1202
Similarity 0.7410 -0.5084 0.1366 0.1240
Table 5: The statistics of Structural bias.
FFHQ 256$\times$256 Iterations
From-scratch Training 80K 90K 100K
DiT-B/2 19.18 18.80 16.55
DiT-B/2 + DMP 15.96 15.21 13.87
Table 6: Applying DMP from the initial training phase.
Appendix H Additional Experiments
Training diffusion models with DMP from scratch.
In Table 6, we show how DMP performs when applied from the start of training. The results indicate that applying DMP from early training stages is still effective, potentially benefiting multi-task learning (MTL) by promoting more nuanced task-specific adaptation via dynamic gating and mixture-of-prompts.
Comparsion with LoRA.
We compare our method with LoRA (Hu et al. 2021) in Table 7. The results show that training a pre-trained diffusion model with LoRA on the original dataset used for pre-training can negatively impact the image generation capability of the diffusion model. In contrast to our DMP, LoRA cannot inherently adjust parameters differently for each timestep. Our DMP, however, enables dynamic control of parameters at each timestep, allowing for more granular adaptation than what is achievable with traditional adapters. This demonstrates that our prompt-based method offers distinct advantages, particularly in terms of ease of implementation and timestep-specific adaptation.
Resolution (256$\times$256) FFHQ
[…]
Pre-trained DiT-B/2 6.27
   + Fine-tuning 6.57(+0.30)
   + Prompt-tuning 6.81(+0.54)
   + LoRA 7.11(+0.84)
   + DMP 5.87(-0.40)
Table 7: Comparison of DMP and LoRA using DiT-B/2 model on FFHQ dataset.
Applying DMP to a large DiT model.
Table 8 shows the results on the DiT-XL/2 model (Peebles and Xie 2022), pre-trained with the ImageNet dataset (Deng et al. 2009). When using the same dataset from the pre-training phase for further training, both fine-tuning and prompt-tuning methods fail to enhance, and even degrade, the class-conditional image generation performance. In contrast, our DMP method effectively improves the model’s performance. In addition, DMP maintains strong precision and recall, matching the base DiT-XL/2 model, while achieving better FID. The improvements were achieved with only 1.26% of additional parameters and a significantly shorter training duration (20K iterations) compared to the full training schedule (7M iterations). This efficiency highlights the practicality and cost-effectiveness of our method, especially for large models where training resources are substantial.
Resolution ($256\times 256$) ImageNet
[…]
Pre-trained (iter: 7M)
DiT-XL/2 2.29 0.83 0.57
Further Training (iter: 20K)
+ Fine-tuning 5.04(+2.75) 0.72 0.61
+ Prompt tuning 2.77(+0.48) 0.81 0.59
+ DMP 2.25(-0.04) 0.83 0.57
Table 8: DMP on DiT-XL/2 model (Peebles and Xie 2022).
Ablation study on gating condition.
In class-conditional image generation and text-to-image generation tasks, conditional guidance plays a crucial role in determining the outcome of generated images (Dhariwal and Nichol 2021; Ho and Salimans 2022). To study the impact of conditional guidance on selecting mixtures-of-prompts, we evaluate the performances of two cases: one where the gating function $\mathcal{G}$ receives only the timestep embedding ${\bm{t}}$, and the other where it receives both ${\bm{t}}$ and the class or text condition embedding ${\bm{c}}$. In the latter case, we modify the input of the gating function in Eq. 3, from ${\bm{t}}$ to ${\bm{t}}+{\bm{c}}$. Our analysis, presented in Tab. 9, indicates that on the ImageNet dataset, both methods perform equivalently, whereas on the COCO dataset, using ${\bm{t}}+{\bm{c}}$ yields superior performance compared to using ${\bm{t}}$ alone. This suggests that incorporating conditional guidance can help in determining how to combine prompts at each denoising step to generate an image that aligns well with the text condition. Consequently, we use ${\bm{t}}+{\bm{c}}$ as the default input for the gating function in the text-to-image task.
@@ Appendix H Additional Experiments
Table 9: Impact of gating condition in DMP.
Additional qualitative results on ImageNet.
@@
Figures 8, 9, 10 and 11 illustrates the generated images by pre-trained DiT-XL with DMP trained on 20K iterations. The results demonstrate that highly realistic images can be generated, even with relatively limited training.
Figure 8: Uncurated 256×256 DiT-XL/2+DMP samples.
@@
Figure 9: Uncurated 256×256 DiT-XL/2+DMP samples.
@@
Figure 10: Uncurated 256×256 DiT-XL/2+DMP samples.
@@
Figure 11: Uncurated 256×256 DiT-XL/2+DMP samples.
@@
Appendix I Limitations
Fixed number of prompts.
Our DMP method adopts a prompt-adding strategy to ensure stable training and maintain sampling speed. However, since the number of input patches is fixed, the flexibility in the number of prompts is limited. Extending our DMP with a prepend approach while maintaining stable training is an interesting future direction.
Different resolutions and aspect ratio.
Our current method, DMP, has limitations in handling different image resolutions and aspect ratios, especially in comparison to models like the FiT (Lu et al. 2024), Lumina-T2X (Gao et al. 2024), or Any-size-diffusion (Zheng et al. 2024), which aim to generalize across arbitrary resolutions and aspect ratios. Unlike these prior works, our DMP requires retraining with different sized prompts specifically tailored to the target resolution and aspect ratio to achieve effective adaptation. While it is possible to use pooling or interpolation techniques to reshape the prompt size for different resolutions and aspect ratios using a pretrained prompt, this approach is arbitrary and can lead to limited performance. Therefore, developing a more flexible method for adapting DMP to varying resolutions and aspect ratios remains an area as further exploration and improvement.